Find the value or values of n in the quadratic equation 2n^2-7n+6=0
Answer by checkley71(8403)

2N~2-7N+6=0 OR (2N-3)(N-2)=0 THEREFORE 2N-3=0 OR 2N=3 OR N=3/2 AND
N-2=0 OR N=2
PROOF 2(3/2)(3/2)-7(3/2)+6=0 OR 18/4-21/2+6=0 OR 18/4-42/4+6=0 OR -24/4+6=0
OR -6+6=0 AND 2(2)(2)-7(2)+6=0 OR 8-14+6=0 OR -6+6=0